# concert_supp

### v2

@@ 4.1 Oracle theoretical results with known $T^{*}$
for some constant $\tilde{c}_{1}>0$.
Theorem 1 can be regarded as an extension of existing results for single-task Bayesian high-dimensional regression (Narisetty and He, 2014; Yang et al., 2016) to the domain of transfer learning. We relax the conditions for single-task Bayesian high-dimensional regression from two perspectives.
First, compared to the requirements outlined in Yang et al. (2016) and Narisetty et al. (2019), where a minimum signal strength constraint of $\min_{j\in S^{*}}(\beta^{(0)*}_{j})^{2}\gtrsim\log p/n_{0}$ is necessary, we relax it in Assumption 4 by replacing the target-only sample size $n_{0}$ in the denominator with a covariate-$j$-specific transferable size $n_{S,T^{*}}^{(j)}$ of all the sources and target. As the number of similar sources along the $j$-th covariate becomes larger, we can detect $\beta_{j}^{(0)*}$ more effectively. In the ideal scenario of $j\in\cap_{k=1}^{K}T_{k}^{*\mathrm{c}}$, we have $n_{S^{*},T^{*}}^{(j)}=N$, where sources are fully utilized. In the worst case where all sources are non-transferable on $S$, i.e., $j\in\cap_{k=1}^{K}T_{k}$, $n_{S^{*},T^{*}}^{(j)}$ remains at $n_{0}$, resembling the case where only target data is utilized.
Second, we allow the dimension $p$ of the covariates to be higher compared to the restriction $\max\{1,s^{*}\}\log p\lesssim n_{0}$ required in (Yang et al., 2016). When all sources are mostly informative, we obtain $\min_{\{S:|S|\leq L_{s}\}}n_{S}=N$, which allows $\log p$ to go as large as the total sample size. Even in the worst case with no informative sources, $\min_{\{S:|S|\leq L_{s}\}}n_{S}$ remains at $n_{0}$, showing the robustness of our method.
Based on the theoretical results on model selection in Theorem 1, we can establish the posterior contraction rates of our proposed method in $\ell_{1}$ loss and $\ell_{2}$ loss to guarantee parameter estimation.
@@
Under the assumptions in Theorem 1 and Assumption LABEL:assump:eigen in the supplementary material, the posterior satisfies that,
@@
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\geq M\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
@@
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\geq M\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
for some constant $\tilde{c}_{2},M>0$.
@@
Theorem 2 indicates that the posterior allocates most of its mass in a neighborhood around the true parameter in terms of $\ell_{1}$ and $\ell_{2}$ loss. The size of the neighborhood depends on $n_{S^{*}}$ indicating that the contraction rates are faster than those when only target data is utilized, thus leading to better estimation behaviors. The $\ell_{1}$ and $\ell_{2}$ error convergences of our method hold under weaker conditions of $(s^{*})^{2}\log p=o(n_{S^{*}})$ and $s^{*}\log p=o(n_{S^{*}})$ than the conventional single-task conditions of $(s^{*})^{2}\log p=o(n_{0})$ and $s^{*}\log p=o(n_{0})$, respectively. Compared with the existing theoretical results of transfer learning (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023), we are able to obtain better error bounds when global similarity is not satisfied.
We can further prove the posterior contraction rate of our proposed method in terms of prediction error under regularity assumptions.
Theorem 3 (Prediction).
Under the assumptions in Theorem 1 and Assumption LABEL:assump:eigen-LABEL:assump:pred in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|X^{(0)}(\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*})\|_{2}\geq M_{1}\sigma^{(0)}_{y}s^{*}\sqrt{\frac{n_{0}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{3}p^{-1},$
for some constant $\tilde{c}_{3},M_{1}>0$.
Similar to Theorem 2, Theorem 3 indicates most of the posterior probability concentrates in the ball around the truth in terms of the prediction error with a faster contraction rate than the single-task Bayesian regression (Castillo et al., 2015).
The main difference between Theorems 1-3 and existing theories of transfer learning (Li et al., 2022; Tian and Feng, 2023; Li et al., 2023) lies in the fact that the benefits of transfer learning brought by our estimator are through each variable rather than the entire set of parameters. Therefore, when there is significant heterogeneity in the source datasets, especially when some variables differ significantly from those in the target dataset, existing methods may fail to extract useful information from the sources due to large overall parameter differences. In contrast, our method can still extract information from similar parts of heterogeneous sources and integrate partial information from each source data to jointly improve the estimation of the target parameters.
4.2 Adaptive theoretical results with unknown $T^{*}$
In practice, the non-transferable sets $\{T^{*}_{k}\}_{k=1}^{K}$ are usually unknown. Therefore, more restrictive assumptions are needed for the simultaneous detection of the significant target signal set and the source-specific non-transferable sets. We provide the following strengthened assumptions.
Assumption 2S (Design matrix).
(i) (Orthogonal design) For $k\in[K]\cup\{0\}$, $\|{{X}}^{(k)}_{j}\|_{2}^{2}=n_{k}$ for all $j\in[p]$ and ${{X}}^{(k)\top}_{j_{1}}{{X}}^{(k)}_{j_{2}}=0$ for all $j_{1}\neq j_{2},j_{1},j_{2}\in[p]$;
(ii) (Restricted eigenvalue) The gram matrix of the considered model space satisfies that $\min_{\{S,T:|S|\leq L_{s},\max_{k}|T_{k}|\leq L_{t}\}}\lambda_{\min}({\tilde{\boldsymbol{X}}_{H_{S},S}(T)^{\top}\tilde{\boldsymbol{X}}_{H_{S},S}(T)}/{n_{S,T}})\geq\lambda_{l}$ for some $\lambda_{l}\in(0,1)$;
$\mathbb{E}_{Z}\left[\max_{S_{1},S_{2}:|S_{1}|\leq|S_{2}|\leq L_{s}}\max_{j\in S_{2}\setminus S_{1}}\max_{T:\max_{k}|T_{k}|\leq L_{t}}\frac{1}{\sqrt{n_{S_{2}}^{(j)}}}\left|{Z}^{T}\tilde{Q}(I_{N}-\Phi_{S_{1},T})\tilde{\boldsymbol{X}}_{j}(T)\right|\right]\leq\frac{\sqrt{C_{3}\lambda_{l}\log p}}{2};$
(iv) (Sparse projection on source $k$) Let ${Z}_{k}\sim\mathcal{N}(0,{I}_{n_{k}})$, $k\in[K]$. For some $C_{3}^{(k)}\geq 8{\lambda_{l}}^{-1}$,
$\mathbb{E}_{Z}\left[\max_{T_{k},\tilde{T}_{k}:|T_{k}|\leq|\tilde{T}_{k}|\leq L_{t}}\max_{j\in\tilde{T}_{k}\setminus{T}_{k}}\left|{Z}_{k}^{T}\tilde{X}^{(k)}_{j}(T_{k})\right|\right]\leq{\sqrt{C_{3}^{(k)}\lambda_{l}\log p}}/{2}.$
Assumption 2S modifies Assumption 2 to account for the case with unknown $T^{*}$. We adopt the orthogonal design Assumption 2S(i) here for simplicity in notation and theoretical analysis. This assumption could be relaxed to a weak correlation assumption, similar to Assumption 2(iv), but with a more refined analysis. Assumption 2S(ii) implicitly considers the model space $\{S,T:\min_{j\in[p]}n_{S,T}^{(j)}\geq c^{\prime}_{0}n_{S,T},|S|\leq L_{s},\max_{k\in[K]}|T_{k}|\leq L_{t}\}$ for some constant $c^{\prime}_{0}>0$. Assumptions 2S(iii) and 2S(iv) are mild and hold automatically with $C_{3}=C_{3}^{(k)}=4\lambda_{l}^{-1}$ under the orthogonal design.
Assumption 3S (Prior).
Assume that the prior satisfies (i) Assumption 3; (ii) $q_{k}\leq p^{-c_{k}}$ with $q_{k}\leq p^{-c_{k}}$ with $c_{k}>2+C_{3}^{(k)}+{3}C_{1}{\lambda_{l}}^{-1}{\underline{\tau}}^{-2}/{2}$ for $k\in[K]$.
@@ 4.2 Adaptive theoretical results with unknown $T^{*}$
In addition to Assumption 3, Assumption 3S(ii) imposes additional assumptions on the prior for the similarity of each source to the target.
Assumption 4S (Large signal).
For all candidate models $T$ satisfying Assumption 2S, assume that the signal strength of covariate $j$ in the target significant set $S^{*}$ satisfies
$\zeta_{j}^{2}\geq C_{\beta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{{n}^{(j)}_{S^{*},T}},$
with constant $C_{\beta}=8({3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/{2}+r+c+2)+4(C_{2}+C_{3})$.
Assumption 4S replaces the ideal sample size $n^{(j)}_{S^{*},T^{*}}$ in the denominator of Assumption 4 with an adaptive ${n}^{(j)}_{S^{*},T}$, which can be seen as a cost of searching for the unknown non-transferable set $T^{*}$. It is important to note that Assumption 4S is no more restrictive than the minimal signal strength of at least the order $\sqrt{\log p/n_{0}}$ needed when using the target data only. This is because $n_{S^{*},T}^{(j)}\geq n_{0}$ always holds, regardless of the model $T$ we consider.
Assumption 5.
Denote $\tilde{n}^{(j)}_{S^{*},T^{*}\setminus T}=\sum_{\{k\in[K]:j\in S^{*}\cap(T_{k}^{*}\setminus T_{k})\}}n_{k}$. Assume that
$\min_{T:\sum_{k}|T_{k}^{*}\setminus T_{k}|\geq 1,\max_{k}|T_{k}|\leq L_{t}}\min_{j\in S^{*}\cap(\cup_{k=1}^{K}T_{k}^{*}\setminus T_{k})}\left(1-\frac{\tilde{n}^{(j)}_{S^{*},T^{*}\setminus T}}{n^{(j)}_{S^{*},T}}\right)\geq C_{l}.$
@@
$\min_{j\in T_{k}^{*}}(\delta_{j}^{(k)*})^{2}\geq C_{\delta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n_{k}},$
for some $C_{\delta}>2C_{l}^{-1}(c_{k}+3C_{1}{{\lambda}_{l}}^{-1}{\underline{\tau}}^{-2}/2+C_{2}/2)+2$.
Assumption 5 and 6 ensure that the probability of missing elements of true model $T^{*}$ tends to zero. Note that in Assumption 5, $n^{(j)}_{S^{*},T}$ is always strictly greater than $\tilde{n}_{S,T^{*}\setminus T}^{(j)}$ as the target set is always included in $H_{S^{*},T}^{(j)}$. Assumption 6 requires the minimal difference strength on the non-transferable signals of source $k$ to be at least the order of $\sqrt{n_{k}/\log p}$ to ensure source dissimilarity detection consistency, which is in the same spirit of the minimal signal strength restriction in Assumption 4S for target signal detection consistency.
Theorem 4 (Model selection).
Suppose $25\max\{1,s^{*}\}\log p\leq\min_{\{S:|S|\leq L_{s}\}}n_{S,T^{*}}$. Under Assumptions 1, 2S-4S, 5, and 6, the posterior of $S^{*},T^{*}$ satisfies that,
@@
for some constant $\tilde{c}_{1}>0$. Consequently, we have the covariate-specific source selection consistency, that is,
$\Pi(S^{*},\{H_{S^{*},T^{*}}^{(j)}\}_{j=1}^{p}|\{y^{(k)}\}_{k=0}^{K})\leq 1-\tilde{c}_{1}p^{-1}.$
Theorem 4 establishes both the target signal detection consistency and the covariate-specific source detection consistency simultaneously, representing a robust and adaptive version of the oracle results in Theorem 1. Regarding robustness, the CONCERT method performs at least as well as the target-only method. For adaptivity, note that the model space of $T$ we search is much smaller in practice. When we restrict our interest to the model space $\mathcal{M}_{T}:=\{T:\min_{j\in[p]}n_{[p],T}^{(j)}\geq c^{\prime}_{0}n_{0}\}$ for some $c^{\prime}_{0}>1$, we can still relax the minimal signal strength requirement in Assumption 4S and detect weaker signals.
Theorem 5 (Estimation).
Under the assumptions in Theorem 4 and Assumption LABEL:assump:eigen in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\geq M\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\geq M\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$

### v3

@@ 4.1.1 Oracle theoretical results with known $T^{*}$
with probability at least $1-\tilde{c}_{3}p^{-\tilde{c}_{4}}$, for some constant $\tilde{c}_{1},\tilde{c}_{3},\tilde{c}_{4}>0$, where the probability is with respect to the data-generating process.
Theorem 1 can be regarded as an extension of existing results for a single-task Bayesian high-dimensional regression (narisetty2014bayesian; yang2016on) to the domain of transfer learning. The proposed transfer learning framework relaxes the conditions required in the single-task setting from two perspectives.
First, compared to the requirements outlined in narisetty2014bayesian and yang2016on, where a minimum signal strength constraint of $\min_{j\in S^{*}}(\beta^{(0)*}_{j})^{2}\gtrsim\log p/n_{0}$ is necessary, we relax it in Assumption 4 by replacing the target-only sample size $n_{0}$ in the denominator with a covariate-$j$-specific informative size $n_{S,T^{*}}^{(j)}$ of all the sources and target. As the number of similar sources along the $j$th covariate becomes larger, we can detect $\beta_{j}^{(0)*}$ more effectively. In the ideal scenario of $j\in\cap_{k=1}^{K}T_{k}^{*\mathrm{c}}$, we have $n_{S^{*},T^{*}}^{(j)}=N$, where all sources are utilized. In the worst case where all sources are non-informative on $S$, i.e., $j\in\cap_{k=1}^{K}T_{k}$, $n_{S^{*},T^{*}}^{(j)}$ remains at $n_{0}$, resembling the case where only target data is utilized.
Second, we allow the dimension $p$ of the covariates to be higher compared to the restriction $\max\{1,s^{*}\}\log p\lesssim n_{0}$ required in yang2016on. When all sources are least-transferable, we obtain $\min_{\{S:|S|\leq L_{s}\}}n_{S}=N$, which allows $\log p$ to go as large as the total sample size. Even in the worst case with no transferable sources, $\min_{\{S:|S|\leq L_{s}\}}n_{S}$ remains at $n_{0}$, showing the robustness of our method.
Based on the model selection results in Theorem 1, we can establish the posterior contraction rates of our proposed method in $\ell_{1}$$\ell_{2}$ loss to guarantee parameter estimation.
@@
Under the assumptions in Theorem 1 and Assumption A.0.1 in the Supplementary Material, the posterior satisfies that,
@@
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\geq M\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*},T^{*}}}}~\bigg|~\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$ (5)
@@
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\geq M\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*},T^{*}}}}~\bigg|~\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{2}p^{-1},$ (6)
with probability at least $1-\tilde{c}_{3}p^{-\tilde{c}_{4}}$, for some constant $\tilde{c}_{2},\tilde{c}_{3},\tilde{c}_{4},M>0$.
@@
Theorem 2 indicates that the posterior allocates most of its mass in a neighborhood around the true parameter in terms of $\ell_{1}$ and $\ell_{2}$ loss. The size of the neighborhood depends on $n_{S^{*}}$, indicating that the contraction rates are faster than those when only target data is utilized, thus leading to better estimation behaviors. The $\ell_{1}$ and $\ell_{2}$ errors converge under weaker conditions of $(s^{*})^{2}\log p=o(n_{S^{*}})$ and $s^{*}\log p=o(n_{S^{*}})$ than the conventional single-task conditions of $(s^{*})^{2}\log p=o(n_{0})$ and $s^{*}\log p=o(n_{0})$, respectively.
In existing transfer learning methods based on global similarity (li2022trans; tian2022transfer), the theoretical results require $\|\boldsymbol{\delta}^{(k)*}\|_{1}\leq s^{*}\sqrt{\log p/n_{0}}$ to ensure positive transfer. As a result, source datasets with $\|\boldsymbol{\delta}^{(k)*}\|_{1}>s^{*}\sqrt{\log p/n_{0}}$ are often excluded from the transferable set. In contrast, by enabling covariate-specific information transfer, our framework allows $\boldsymbol{\delta}^{(k)*}$ to be large within $T_{k}^{*}$, thereby achieving improved error bounds when the global similarity requirement on $\|\boldsymbol{\delta}^{(k)*}\|_{1}$ is not satisfied.
4.1.2 Adaptive theoretical results with unknown $T^{*}$
In practice, the non-transferable sets $\{T^{*}_{k}\}_{k=1}^{K}$ are usually unknown. This necessitates the simultaneous detection of the significant target signal set and the source-specific non-transferable sets, which is more challenging. To achieve this, we make slight adjustments to Assumptions 2 and 4 to accommodate the case where $T$ is unknown. Due to space limitations, we present them as Assumptions B.2.1 and B.2.2 in Section B.2.1 of the Supplementary Material. Let $C_{3}^{(k)}$ be some large number introduced in Assumption B.2.1. Here we present the following additional assumptions.
Assumption 5 (Prior).
The prior satisfies (i) $q_{k}\sim p^{-1}$; (ii) $n_{k}\tau_{k}^{2}\sim(n_{k}\vee p^{2r_{k}})$ for some $r_{k}>1+r+3/2C_{1}\lambda_{l}^{-1}+C_{3}^{(k)}$ for $k\in[K]$.
@@ 4.1.2 Adaptive theoretical results with unknown $T^{*}$
In addition to Assumption 3, Assumption 5 imposes assumptions on the prior for the similarity of each source to the target.
@@
$\min_{j\in T_{k}^{*}}(h_{j}^{(k)})^{2}\geq C_{\delta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n_{k}d_{1}d_{2}},$
for some $C_{\delta}=8(3C_{1}{{\lambda}_{l}}^{-1}/2+C_{2}/2+r_{k}+3+(C_{2}+C_{3}^{(k)})/4)$.
Assumption 6 requires the minimal heterogeneity level on the non-transferable signals of source $k$ to be at least the order of $\sqrt{\log p/n_{k}}$ to ensure source dissimilarity detection consistency, which is in the same spirit as the minimal signal strength restriction in Assumption 4 for target signal detection consistency.
Theorem 3.
Suppose $\min_{\{S:1\leq|S|\leq L_{s}\}}n_{S,T^{*}}\gtrsim\log p$. Under Assumptions 1, 3, 5, 6, and B.2.1-B.2.2, the posterior of $S^{*},T^{*}$ satisfies that,
@@
with probability at least $1-\tilde{c}_{3}p^{-\tilde{c}_{4}}$, for some constant $\tilde{c}_{1},\tilde{c}_{3},\tilde{c}_{4}>0$. Consequently, we have the covariate-specific source selection consistency, that is,
$\Pi(S^{*},\{H_{S^{*},T^{*}}^{(j)}\}_{j=1}^{p}|\{y^{(k)}\}_{k=0}^{K})\geq 1-\tilde{c}_{1}p^{-1},$
with probability at least $1-\tilde{c}_{3}p^{-\tilde{c}_{4}}$.
Furthermore, under Assumption A.0.1, the posterior contraction rates in terms of $\ell_{1}$ loss and $\ell_{2}$ loss in (5)-(6) holds.
The first part of Theorem 3 establishes both the target signal detection consistency and the covariate-specific source detection consistency simultaneously. The second part of Theorem 3 demonstrates that the CONCERT achieves desirable posterior contraction rates in terms of $\ell_{1}/\ell_{2}$ loss under the unknown $T^{*}$ case. We remark that Theorem 3 is a non-trivial extension of the oracle results in Section 4.1.1 due to the interplay between variable selection and transferability assessment. To overcome this, we use a refined decomposition of the model space to enable a more precise and structured theoretical analysis.
4.2 VB Posterior Contraction
In this subsection, we present the contraction rates for the VB posterior. Building upon the contraction results established for the true posterior, we now directly consider the case where $T$ is unknown. To this end, we refine our assumptions regarding the prior and signal strength, as detailed in Assumptions B.3.1-B.3.3 in Section B.3.1 of the Supplementary Material. Let $\{\kappa_{n}\}$ be a sequence that tends to infinity at an arbitrarily slow rate, as introduced in Assumption B.3.1. We have the following results.
Theorem 4 (Model selection under VB posterior).
Suppose $\min_{\{S:1\leq|S|\leq L_{s}\}}n_{S,T^{*}}\gtrsim\log p$. Under Assumptions 1, B.2.1, and B.3.1-B.3.3, the VB posterior satisfies that,
$\mathbb{E}_{\boldsymbol{\theta}^{*}}Q^{*}(S^{*},T^{*})\geq 1-O(1/\kappa_{n})-\tilde{c}_{1}p^{-1},$
for some constant $\tilde{c}_{1}>0$, where $\mathbb{E}_{\boldsymbol{\theta}^{*}}$ denotes the expectation with respect to the probability measure of the true data-generating process under $\boldsymbol{\theta}^{*}$.
Compared to the corresponding result for the true posterior in Theorem 3, Theorem 4 includes an additional $O(1/\kappa_{n})$ term due to the variational approximation, implying that the VB posterior concentrates on the true model at a slightly slower rate. A similar $O(1/\kappa_{n})$ term also appears in the theoretical analysis of VB by ray2022vb.
Similarly, we can establish the VB posterior contraction rates in terms of $\ell_{1}$ and $\ell_{2}$ loss.
Theorem 5 (Estimation under VB posterior).
Under the assumptions in Theorem 4 and Assumption A.0.1, the VB posterior satisfies that,
$\mathbb{E}_{\boldsymbol{\theta}^{*}}Q^{*}\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{1}\leq M\kappa_{n}\sigma^{(0)}_{y}s^{*}\sqrt{\frac{\log p}{n_{S^{*},T^{*}}}}\right)\geq 1-O(1/\kappa_{n})-\tilde{c}_{2}p^{-1},$
$\mathbb{E}_{\boldsymbol{\theta}^{*}}Q^{*}\left(\boldsymbol{\beta}^{(0)}:\|\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*}\|_{2}\leq M\kappa_{n}\sigma^{(0)}_{y}\sqrt{\frac{s^{*}\log p}{n_{S^{*},T^{*}}}}\right)\geq 1-O(1/\kappa_{n})-\tilde{c}_{2}p^{-1},$